# Bayesian Controlled FDR Variable Selection via Knockoffs

### v1

@@ Abstract
In many research fields, researchers aim to identify significant associations between a set of explanatory variables and a response while controlling the false discovery rate (FDR). To this aim, we develop a fully Bayesian generalization of the classical model-X knockoff filter. Knockoff filter introduces controlled noise in the model in the form of cleverly constructed copies of the predictors as auxiliary variables. In our approach we consider the joint model of the covariates and the response and incorporate the conditional independence structure of the covariates into the prior distribution of the auxiliary knockoff variables. We further incorporate the estimation of a graphical model among the covariates, which in turn aids knockoffs generation and improves the estimation of the covariate effects on the response. We use a modified spike-and-slab prior on the regression coefficients, which avoids the increase of the model dimension as typical in the classical knockoff filter. Our model performs variable selection using an upper bound on the posterior probability of non-inclusion. We show how our model construction leads to valid model-X knockoffs and demonstrate that the proposed characterization is sufficient for controlling the BFDR at an arbitrary level, in finite samples. We also show that the model selection is robust to the estimation of the precision matrix. We use simulated data to demonstrate that our proposal increases the stability of the selection with respect to classical knockoff methods, as it relies on the entire posterior distribution of the knockoff variables instead of a single sample. With respect to Bayesian variable selection methods, we show that our selection procedure achieves comparable or better performances, while maintaining control over the FDR. Finally, we show the usefulness of the proposed model with an application to real data.
Keywords: False Discovery Rate; Graphical Models; MCMC; Regression; Spike-and-Slab Priors
@@ 1 Introduction
In recent years, variable selection has emerged as a crucial task at the forefront of statistics and machine learning. Variable selection is becoming increasingly essential in scientific research, as it enables to search for associations between an outcome variable and a large amount of possible explanatory variables, despite only a few of them being truly relevant. Thus, the need for methods that can accurately identify the subset of variables that affect the outcome is paramount. Consequently, variable selection is a well-studied topic in the statistical literature, particularly in the regression context, where it reduces to the selection of the non-zero model coefficients. Classical approaches for variable selection typically rely on constrained optimization of the likelihood function (see, among others, Tibshirani,, 1996; Fan and Li,, 2001; Zou and Hastie,, 2005). In the Bayesian approach, variable selection can be performed by employing different classes of prior distributions (Tadesse and Vannucci,, 2021). For regression models, spike-and-slab priors, proposed for the first time by Mitchell and Beauchamp, (1988), impose mixture distributions on the regression coefficients, with latent binary indicators that identify the active predictors (George and McCulloch,, 1997; Brown et al.,, 1998). Another approach involves the use of shrinkage priors to adaptively shrink coefficients towards zero (Park and Casella,, 2008; Carvalho et al.,, 2010).
@@
A fundamental challenge in variable selection is controlling the False Discovery Rate (FDR), to ensure the reliability and replicability of research findings. False discoveries refer to situations where a statistical procedure incorrectly identifies a variable as active when, in reality, its effect is merely due to random chance or noise in the data. Multiple testing procedures have been proposed (see, for instance, Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001, 2005) that aim at controlling the average proportion of false discoveries. These procedures typically rely on computing $p$-values, which is not always an easy task (Candès et al.,, 2018). Alternative methodologies have been developed by Xing et al., (2023), who propose a method that leverages a pair of mirror variables for each predictor to provide asymptotic control on the FDR, and by Dai et al., (2022), who employ a data-splitting procedure to accomplish the same goal with minimal computational effort.
Certain classical techniques for FDR control reach their goal by introducing controlled perturbation to the data. Barber and Candès, (2015) propose the knockoff filter, a powerful tool that effectively achieves finite sample FDR control in fixed design low dimensional linear models. The key idea is to add controlled noise in the form of cleverly constructed copies of the predictors. Subsequently, a standard linear model is fitted on the augmented design matrix, and variable selection is performed using a function of the coefficients of both the original and the knockoff variables. Basically, knockoff variables are used as negative controls. This idea is extended to the case of fixed design high dimensional linear models in Barber and Candès, (2019). Since this approach relies on a fixed design matrix, it is called fixed-X knockoff. An extension of this setting is the so-called model-X knockoff filter (Candès et al.,, 2018), which does not assume as true any model on the relation between the response and the covariates. Instead, it requires a complete knowledge of the joint distribution of the covariates. Alternative perspectives have been investigated by Sesia et al., (2018), who provide an efficient exact sampling of the knockoff variables when the covariates are distributed as a hidden Markov model, and by Bates et al., (2021), who apply a modified version of the algorithm proposed by Candès et al., (2018) to sample knockoffs exactly in the case of known independence structure of the covariates. Finally, Berti et al., (2023) provide a new characterization of the joint distribution of the observed covariates and the knockoffs, which they define in terms of copulas. When the covariates follow a joint Gaussian distribution their approach provides the same solution as in Candès et al., (2018).
@@
In Bayesian settings, the usual definition of FDR cannot be used, since both the set of active and inactive variables are considered random variables. Instead, the Bayesian FDR (BFDR) is defined as an expected FDR conditionally upon the observed data, that is as the expected proportion of incorrect rejections among all possible rejections, computed with respect to the posterior distribution of active/inactive variables conditioned upon the observed data (Muller et al.,, 2006; Whittemore,, 2007). An advantage of Bayesian methods is that they can facilitate multiplicity adjustments through the choice of the prior. For example, in the context of variable selection, Scott and Berger, (2010) show that a spike-and-slab prior with a Beta-Bernoulli prior on the selection indicator achieves an asymptotic multiplicity control. Muller et al., (2006) and Guindani et al., (2009) prove control of the Bayesian FDR by characterizing the FDR problem as a decision problem where the optimal decision rule is based on thresholding the marginal posterior probabilities. More recently, Castillo and Roquain, (2020) prove control of the Bayes FDR for general empirical Bayes multiple testing procedures. In practical settings, the optimal threshold for controlling the FDR corresponds to the minimum marginal probability of inclusion that ensures that the BFDR is less than the desired level (Newton et al.,, 2004). Therefore, this procedure relies on the estimates of these probabilities, as obtained from the MCMC output.
@@
In this work, we develop a fully Bayesian generalization of the model-X knockoff filter. One of the main issues of the classical model-X knockoff filter is that the inference process concerning the set of active variables heavily relies on the quality of the randomly generated sample of knockoff variables. As a result, the procedure may exhibit instability and produce different sets of active variables for the same set of predictors. To mitigate this issue, Ren et al., (2023) combine the knockoff filter with the stability selection procedure of Meinshausen and Bühlmann, (2010), obtaining improved inference and a more stable selection of active variables. This issue of stability is naturally addressed within the Bayesian framework, where the knockoffs variables are treated as latent and updated as part of the posterior inference. In particular, in our proposed approach, we consider the joint model of the covariates and the response and incorporate the conditional independence structure of the covariates into the prior distribution of the knockoff variables. Our joint modeling construction employs a likelihood decomposition via auxiliary variables that allows us to obtain valid model-X knockoffs. We further assume a Gaussian graphical model on the covariates and impose the continuous spike-and-slab prior of Wang, (2015) on the precision matrix. Estimating a graphical model on the covariates allows us to obtain a coherent model which does not require perfect knowledge of the precision matrix. Moreover, differently from any plug-in estimator, we can incorporate prior information about the conditional dependence structure of the covariates into the model, assuring a better knockoff quality, as argued by Barber et al., (2020). For the regression coefficients, we follow a suggestion from Candès et al., (2018) and assume a mixture of spike-and-slab prior, which avoids the increase of the model dimension typical with the classical knockoff filter, and which we further modify by letting the probability of inclusion of a variable depend on the number of its neighbors in the graph. We ensure that the knockoff variables possess a priori the required properties outlined in Candès et al., (2018) and demonstrate that this characterization is sufficient for controlling the BFDR at an arbitrary level, in finite samples. Moreover, for a sufficiently large sample size, we prove that FDR control is achieved even without the knowledge of the true covariates’ graph. In applications, to maintain control over the BFDR, we employ a greedy algorithm, similar to the one presented in Newton et al., (2004), but relying on an estimate of the upper bound for the probability of non-inclusion. We use simulated data to demonstrate that our proposal increases the stability of the selection with respect to classical knockoff methods, as it relies on the entire posterior distribution of the knockoff variables instead of a single sample. With respect to Bayesian variable selection methods, we show that our selection procedure achieves comparable or better performances, while maintaining control over the FDR.
@@
Very few attempts have been made at constructing Bayesian knockoffs filters. Gu and Yin, (2021) first considered a generalized linear model framework and assumed the distribution of the covariates to be completely known. With respect to their model formulation, in our approach the incorporation of the estimation of a graphical model on the covariates aids knockoffs generation and improves the estimation of the covariate effects on the response. Furthermore, unlike the model formualtion of Gu and Yin, (2021), our joint model ensures valid model-X knockoffs. Yap and Gauran, (2023) employ the Bayesian lasso of Park and Casella, (2008) for the estimation of the regression coefficients. However, they treat the covariates as fixed and assume that the augmented knockoff matrix is generated before the application of the Bayesian method.
@@ 2 Methods
2.1 Knockoff filter
Knockoff filter is a framework that enables variable selection in an FDR-controlled environment. The main idea, originally proposed by Barber and Candès, (2015), is to perturb the model by adding a set of inactive predictors built in such a way that they resemble the dependence structure of the original covariates. In Barber and Candès, (2015) the knockoff predictors were built deterministically, while Candès et al., (2018) extended the framework to the model-$X$ knockoff filter, with random covariates. The model-$X$ knockoff assumes a joint distribution $p(X)$ for the predictors and random knockoff variables, $\widetilde{X}=(\widetilde{X}_{1},\dots,\widetilde{X}_{p})^{T}$, which are defined in a way to make it hard for the model to distinguish them from $X$. Formally, the knockoffs are sampled from a distribution $p(\widetilde{X}\mid X)$ such that these two properties hold:
@@ 2.1 Knockoff filter
$A=2\text{diag}(s)-\text{diag}(s)\bm{\Sigma}^{-1}\text{diag}(s)$ (2)
@@
with $U\sim N(0,A)$.
2.2 Bayesian knockoffs
@@
$Y=X[\beta+(I_{p}-\bm{\Omega}\text{diag}(s))\widetilde{\beta}]+U\widetilde{\beta}+\epsilon,\qquad\epsilon\sim N(0,\sigma^{2}),$ (8)
@@
with the matrix $A$ as defined in Equation (2). This construction respects the conditional independence property of the knockoff variables, as now $Y\mbox{$\perp\!\!\!\perp\,$}\widetilde{X}\mid X$. Moreover, since the decomposition of the knockoff variables in Equation (4) is obtained from the joint distribution in Equation (1), our formulation respects the pairwise exchangeability too. Therefore, this setting uses valid model-X knockoffs.
@@ 2.2.2 Priors for variable selection
For the regression coefficients, we consider a modified spike-and-slab prior distribution similar to the one introduced in the supplementary material of Candès et al., (2018). This prior is specified as the joint distribution
@@
with $j=1,\ldots,p$. Here, $h_{\beta}$ is a fixed positive scaling hyperparameter and $\gamma=(\gamma_{1},\dots,\gamma_{p})^{T}\in\{0,1\}^{p}$ is a $p$-dimensional vector of binary indicators such that $\gamma_{j}=1$ if $X_{j}$ or its knockoff $\widetilde{X}_{j}$ are included in the model and, alternatively, $\gamma_{j}=0$ if neither the original feature nor the knockoff is included. By introducing two additional latent indicator vectors, $\delta=(\delta_{1},\dots,\delta_{p})$ and $\widetilde{\delta}=(\widetilde{\delta}_{1},\dots,\widetilde{\delta}_{p})$, such that $\gamma=\delta+\widetilde{\delta}$, the joint prior in Equation (12) can be reformulated as a mixture of independent spike-and-slab priors on $\beta_{j}$ and $\tilde{\beta}_{j}$ as
@@ 2.2.3 Markov Chain Monte Carlo for posterior sampling
This distribution is not analytically tractable, and we therefore resort to a Markov Chain Monte Carlo (MCMC) algorithm (Brooks et al.,, 2011). In particular, we implement an efficient Metropolis within Gibbs scheme that employs a data augmentation (DA) scheme based on the latent variable $U$ introduced in (4) and the formulation of the joint model detailed in Section 2.2. Specifically, as it is not possible to obtain tractable full conditional distributions using the marginal model $p(Y\mid X,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2})$, we employ a DA step by sampling from the distribution $p(Y\mid X,U,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2})$ defined in Equation (8) and then updating $U$ from the full conditional. We also incorporate the Add-Delete Metropolis-Hastings algorithm of Savitsky et al., (2011), that cleverly avoids having to deal with the changing dimensions of the parameter space via a joint update of the inclusion indicators of the spike-and-slab priors and the corresponding regression coefficients. In practice, our algorithm is composed by the following steps:
@@
    We jointly update $(\delta,\widetilde{\delta},\gamma,\beta,\widetilde{\beta})$ with a stochastic search approach. At each iteration $t=1,\dots,T$, we randomly select one $j\in\{1,\dots,p\}$ and propose to change the value of $\gamma_{j}$. If the proposed $\gamma_{j}^{*}=0$ then the only possible choice for the proposed $\beta_{j}^{*}$, $\widetilde{\beta}_{j}^{*}$, $\delta_{j}^{*}$, $\widetilde{\delta}_{j}^{*}$ is to set them all to 00. However, if we select a non-included variable, $\gamma_{j}=0$, we randomly propose to include the original variable or its knockoff. More specifically, we propose the selected coefficient using a Gaussian distribution $N(0,0.5)$ as the proposal distribution. We then accept the proposed parameters with probability
    $r=\min\left\{1,\frac{p(Y\mid X,U,\beta^{*},\widetilde{\beta}^{*},\bm{\Omega},\sigma^{2})p(\beta^{*},\widetilde{\beta}^{*}\mid\gamma^{*},\sigma^{2},h_{\beta})p(\delta^{*},\widetilde{\delta}^{*}\mid\gamma^{*})p(\gamma^{*}\mid\bm{G},a,b)}{p(Y\mid X,U,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2})p(\beta,\widetilde{\beta}\mid\gamma,\sigma^{2},h_{\beta})p(\delta,\widetilde{\delta}\mid\gamma)p(\gamma\mid\bm{G},a,b)}\right\}$
@@
    $\sigma^{2}\mid y,\bm{X},U,\beta,\widetilde{\beta},\bm{\Omega}\sim IG\left(a_{\sigma}+\frac{n}{2},b_{\sigma}+\frac{\sum_{i}(y_{i}-x_{i}[\beta+(I-\bm{\Omega}\text{diag}(s))\widetilde{\beta}]-U_{i}\widetilde{\beta})^{2}}{2}\right).$
@@
    $u_{i}\mid y,\bm{X},U,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2}\sim N(\mu_{i},\Sigma_{U})$
    $\Sigma_{U}=\left(A+\frac{1}{\sigma^{2}}\widetilde{\beta}\,\widetilde{\beta}^{T}\right)^{-1}$
    $\mu_{i}=\Sigma_{U}\left(\frac{1}{\sigma^{2}}\left(y_{i}\widetilde{\beta}-\widetilde{\beta}\,\beta^{T}x_{i}-\widetilde{\beta}\widetilde{\beta}^{T}\Gamma_{i}\right)\right)$
    $\Gamma=X(I-\Omega\text{diag}(s)),$

### v2

@@ Abstract
In many research fields, researchers aim to identify significant associations between a set of explanatory variables and a response while controlling the false discovery rate (FDR). The Knockoff filter has been recently proposed in the frequentist paradigm to introduce controlled noise in a model by cleverly constructing copies of the predictors as auxiliary variables. In this paper, we develop a fully Bayesian generalization of the classical model-X knockoff filter for normally distributed covariates. In our approach we consider a joint model of the covariates and the response variables, and incorporate the conditional independence structure of the covariates into the prior distribution of the auxiliary knockoff variables. We further incorporate the estimation of a graphical model among the covariates, leading to improved knockoffs generation and estimation of the covariate effects on the response. We use a modified spike-and-slab prior on the regression coefficients, which avoids the increase of the model dimension as typical in the classical knockoff filter. Our model performs variable selection using an upper bound on the posterior probability of non-inclusion. We show how our construction leads to valid model-X knockoffs and demonstrate that the proposed variable selection procedure leads to controlling the Bayesian FDR at an arbitrary level, in finite samples, if the distribution of the covariates is fully known, and asymptotically if estimated as in the proposed model. We use simulated data to demonstrate that our proposal increases the stability of the selection with respect to classical knockoff methods, as it relies on the entire posterior distribution of the knockoff variables instead of a single sample. With respect to Bayesian variable selection methods, we show that our selection procedure achieves comparable or better performances, while maintaining control over the FDR. Finally, we show the usefulness of the proposed model with an application to real data.
Keywords: False Discovery Rate, Graphical Models, MCMC, Regression, Spike-and-Slab Priors.
@@ 1 Introduction
In recent years, variable selection has emerged as a crucial task at the forefront of statistics and machine learning, since it enables searching for associations between an outcome variable and a large amount of possible explanatory variables, when only a few of them are truly relevant. Hence, there is a need for methods that can accurately identify the subset of variables that affect the outcome. Classical approaches for variable selection typically rely on constrained optimization of the likelihood function (see, among others, Tibshirani,, 1996; Fan and Li,, 2001; Zou and Hastie,, 2005). In the Bayesian approach, variable selection can be performed by employing several classes of prior distributions (Tadesse and Vannucci,, 2021, for a review, see). For regression models, spike-and-slab priors, proposed for the first time by Mitchell and Beauchamp, (1988), impose mixture distributions on the regression coefficients, with latent binary indicators that identify the relevant (active) predictors (George and McCulloch,, 1997; Brown et al.,, 1998). Another approach involves the use of shrinkage priors to adaptively shrink coefficients towards zero (Park and Casella,, 2008; Carvalho et al.,, 2010).
@@
A fundamental challenge in variable selection is controlling the False Discovery Rate (FDR), to ensure the reliability and replicability of research findings. False discoveries refer to situations where a statistical procedure incorrectly identifies a variable as relevant when, in reality, its effect is merely due to random chance or noise in the data. Several multiple testing procedures have been proposed to control the FDR - that is, the expected proportion of inactive variables mistakenly identified as active, among all those selected as active (see, for instance, Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001, 2005). Those procedures typically rely on computing $p$-values, which is not always an easy task (Candès et al.,, 2018). Alternative methodologies have been developed by Xing et al., (2023), who propose a method that leverages a pair of mirror variables for each predictor to provide asymptotic control on the FDR, and by Dai et al., (2022), who employ a data-splitting procedure to accomplish the same goal with minimal computational effort.
@@
In Bayesian settings, the usual definition of FDR cannot be used, since both the set of active and inactive variables are considered random. Instead, the Bayesian FDR (BFDR) is defined as an expected FDR conditionally upon the observed data, i.e., computed with respect to the posterior distribution of the sets of active/inactive variables given the observed data (Muller et al.,, 2006; Whittemore,, 2007). An advantage of Bayesian methods is that they can facilitate multiplicity adjustments through the choice of the prior. For example, in the context of variable selection, Scott and Berger, (2010) show that a spike-and-slab prior with a Beta-Bernoulli prior on the selection indicator achieves an asymptotic multiplicity control. Muller et al., (2006) and Guindani et al., (2009) prove control of the Bayesian FDR by characterizing the FDR problem as a decision problem where the optimal decision rule is based on thresholding the marginal posterior probabilities. More recently, Castillo and Roquain, (2020) prove control of the Bayes FDR for general empirical Bayes multiple testing procedures. In practical settings, the optimal threshold for controlling the FDR corresponds to the minimum marginal probability of inclusion that ensures that the BFDR is less than a desired level (Newton et al.,, 2004). Therefore, this procedure relies on the estimates of these probabilities, which are typically obtained from the MCMC output.
Recent classical techniques to control the FDR do so by introducing carefully designed perturbations to the data. For example, Barber and Candès, (2015) propose the knockoff filter, a powerful method that achieves finite sample FDR control in fixed-design, low-dimensional linear models. The key idea is to add controlled noise in the form of cleverly constructed copies of the predictors. These copies, called knockoffs, are constructed to mimic the correlation structure of the true covariates while being statistically independent of the response. A linear model is then fitted to this augmented design, and variable selection is performed using a function of the coefficients of both the original and the knockoff variables. Essentially, knockoff variables act as negative controls. This framework has been extended to high-dimensional fixed-design settings in Barber and Candès, (2019), and is often referred to as the fixed-X knockoff procedure. The model-X knockoff filter of Candès et al., (2018) further extends the approach, by avoiding assumptions on the relation between the response and the covariates. Instead, it requires full knowledge of the joint distribution of the covariates. Alternative perspectives have been investigated by Sesia et al., (2018), who provide an efficient exact sampling of the knockoff variables when the covariates are distributed as a hidden Markov model, and by Bates et al., (2021), who apply a modified version of the algorithm proposed by Candès et al., (2018) to sample knockoffs exactly in the case of known independence structure of the covariates. More recently, Berti et al., (2023) provide a new characterization of the joint distribution of the observed covariates and the knockoffs, which they define in terms of copulas. When the covariates follow a joint Gaussian distribution, their approach recovers the same solution as in Candès et al., (2018).
One of the main limitations of the classical model-X knockoff filter is that it depends on the quality of a single random draw of knockoff variables: different draws, from the same set of predictors, may yield different selected sets. To reduce this variability, Ren et al., (2023) combine the knockoff filter with the stability-selection framework of Meinshausen and Bühlmann, (2010), yielding a more stable selection of active variables.
In this work, we take a fully Bayesian approach that generalizes the model-X knockoff filter and builds stability directly into the inference process. More specifically, we treat the knockoffs as latent variables and update them as part of the posterior inference. We consider the joint model of the covariates and the response, and incorporate the conditional independence structure of the covariates into the prior distribution of the knockoff variables. Our joint modeling construction employs a likelihood decomposition via auxiliary variables that allows us to obtain valid model-X knockoffs.
We further assume a Gaussian graphical model to model the covariates’ dependence structure and impose a continuous spike-and-slab prior on the precision matrix (Wang,, 2015). Estimating a graphical model on the covariates allows us to obtain a coherent model that does not require perfect knowledge of the precision matrix. Moreover, unlike plug-in estimators, our Bayesian approach can naturally incorporate prior information about the conditional dependence structure of the covariates into the model. As argued by Barber et al., (2020), this leads to higher-quality knockoffs.
@@
For the regression coefficients, we adopt the mixture spike-and-slab prior suggested by Candès et al., (2018), which avoids the increase of the model dimension typical of the classical knockoff filter. Additionally, we modify this prior by allowing the probability of inclusion of each variable to depend on the number of its neighbors in the graph. We ensure that the knockoff variables possess a priori the required properties outlined in Candès et al., (2018) and demonstrate that this characterization is sufficient for controlling the BFDR at an arbitrary level, in finite samples, provided the distribution of the covariates is fully known. Moreover, for a sufficiently large sample size, our procedure maintains asymptotic FDR control when the true covariates’ graph is estimated from data, as in our proposed framework. In applications, to maintain control over the BFDR, we employ a greedy algorithm, similar to the one presented in Newton et al., (2004), but relying on an estimate of the upper bound for the probability of non-inclusion. Using simulated data, we show that our proposal substantially improves the stability of the selection compared to classical knockoff methods, as it leverages the entire posterior distribution of the knockoff variables rather than relying on a single realization. Furthermore, we show that our procedure achieves performance comparable to or better than standard Bayesian variable selection methods, while consistently controlling the FDR.
@@
To our knowledge, only a few Bayesian knockoff filters have been proposed in the literature to date. Gu and Yin, (2021) first considered a generalized linear model framework and assumed the distribution of the covariates to be completely known. With respect to their model formulation, in our approach, the incorporation of the estimation of a graphical model on the covariates aids in knockoff generation and improves the estimation of the covariate effects on the response. Furthermore, unlike the model formulation of Gu and Yin, (2021), our joint model ensures valid model-X knockoffs. Yap and Gauran, (2023) employ the Bayesian lasso of Park and Casella, (2008) for the estimation of the regression coefficients. However, they treat the covariates as fixed and require the augmented knockoff matrix to be generated prior to Bayesian inference.
@@ 2 Methods
2.1 A review of the Knockoff filter
Knockoff filter is a framework that enables variable selection in an FDR-controlled environment. The main idea, originally proposed by Barber and Candès, (2015), is to perturb the model by adding a set of inactive predictors built in such a way that they resemble the dependence structure of the original covariates. In Barber and Candès, (2015) the knockoff predictors were built deterministically, while Candès et al., (2018) extended the framework to the model-$X$ knockoff filter, with random covariates. The model-$X$ knockoff assumes a joint distribution $p(X)$ for the predictors and random knockoff variables, $\widetilde{X}=(\widetilde{X}_{1},\dots,\widetilde{X}_{p})$, which are defined in a way to make it hard for the model to distinguish them from $X$. Formally, the knockoffs are sampled from a distribution $p(\widetilde{X}\mid X)$ such that these two properties hold:
@@ 2.1 A review of the Knockoff filter
${\bm{A}}=2\text{diag}(s)-\text{diag}(s)\bm{\Sigma}^{-1}\text{diag}(s)$ (2)
@@
with $U\sim N(0,\bm{A})$.
2.2 Constructing Bayesian knockoffs
@@
$Y=X\left[\beta+(I_{p}-\bm{\Omega}\text{diag}(s))\widetilde{\beta}\right]+U\widetilde{\beta}+\epsilon,\qquad\epsilon\sim N(0,\sigma^{2}),$ (8)
@@
with the matrix ${\bm{A}}$ as defined in Equation (2). This construction respects the conditional independence property of the knockoff variables, as now $Y\mbox{$\perp\!\!\!\perp\,$}\widetilde{X}\mid X$. Moreover, since the decomposition of the knockoff variables in Equation (4) is obtained from the joint distribution in Equation (1), our formulation respects the pairwise exchangeability too. Therefore, this setting uses valid model-X knockoffs. Note that the construction of the Bayesian knockoff filter proposed by Gu and Yin, (2021) does not produce valid knockoffs, since adopting $p(Y\mid X,\widetilde{X})$ as the likelihood function, as they do, implies $Y\mbox{$\perp\!\!\!\perp\!\!\!\!\!\!\setminus\;$}\widetilde{X},X$ and therefore the conditional independence property is not satisfied.
@@ 2.2.2 Priors for variable selection
For the regression coefficients, we adapt the spike-and-slab prior introduced in the supplementary material of Candès et al., (2018). Specifically, we modify their formulation by allowing each variable’s inclusion probability to depend on its number of neighbors in the estimated covariate graph, as detailed below. For simplicity of exposition, here we refer to the conditional model in (7), although similar arguments apply to the marginal model (9). We first specify the joint distribution
@@
with $j=1,\ldots,p$. Here, $h_{\beta}$ is a positive scaling hyperparameter and $\gamma=(\gamma_{1},\dots,\gamma_{p})^{T}\in\{0,1\}^{p}$ is a $p$-dimensional vector of binary indicators such that $\gamma_{j}=1$ if $X_{j}$ or its knockoff $\widetilde{X}_{j}$ are included in the model and, alternatively, $\gamma_{j}=0$ if neither the original feature nor the knockoff is included. By introducing two additional latent indicator vectors, $\delta=(\delta_{1},\dots,\delta_{p})$ and $\widetilde{\delta}=(\widetilde{\delta}_{1},\dots,\widetilde{\delta}_{p})$, such that $\gamma=\delta+\widetilde{\delta}$, the joint prior in Equation (12) can be reformulated as a mixture of independent spike-and-slab priors on $\beta_{j}$ and $\tilde{\beta}_{j}$ as
@@ 2.2.3 Markov Chain Monte Carlo for posterior sampling
This distribution is not analytically tractable, and we therefore resort to a Markov Chain Monte Carlo (MCMC) algorithm (Brooks et al.,, 2011). In particular, we implement an efficient Metropolis within Gibbs scheme that employs a data augmentation (DA) scheme based on the latent variable $U$ introduced in (4) and the formulation of the joint model detailed in Section 2.2. Specifically, as it is not possible to obtain tractable full conditional distributions using the marginal model $p(Y\mid X,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2})$, we employ a DA step by sampling from the distribution $p(Y\mid X,U,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2})$ defined in Equation (8) and then updating $U$ from the full conditional. We also incorporate the Add-Delete Metropolis-Hastings algorithm of Savitsky et al., (2011), that cleverly avoids having to deal with the changing dimensions of the parameter space via a joint update of the inclusion indicators of the spike-and-slab priors and the corresponding regression coefficients. We provide full details of the MCMC algorithm in the Supplementary Material. Briefly, our algorithm is composed by the following steps:
@@
    We jointly update $(\delta,\widetilde{\delta},\gamma,\beta,\widetilde{\beta})$ with a stochastic search approach. At each iteration $t=1,\dots,T$, we randomly select one $j\in\{1,\dots,p\}$ and propose to change the value of $\gamma_{j}$. If the proposed $\gamma_{j}^{*}=0$ then the only possible choice for the proposed $\beta_{j}^{*}$, $\widetilde{\beta}_{j}^{*}$, $\delta_{j}^{*}$, $\widetilde{\delta}_{j}^{*}$ is to set them all to 00. However, if we select a non-included variable, $\gamma_{j}=0$, we randomly propose to include the original variable or its knockoff. More specifically, we propose the selected coefficient using a Gaussian distribution $N(0,0.5)$ as the proposal distribution. We then accept the proposed parameters with probability $r=\min\left\{1,\rho\right\}$ with
    $\rho=\frac{p(Y\mid X,U,\beta^{*},\widetilde{\beta}^{*},\bm{\Omega},\sigma^{2})p(\beta^{*},\widetilde{\beta}^{*}\mid\gamma^{*},\sigma^{2},h_{\beta})p(\delta^{*},\widetilde{\delta}^{*}\mid\gamma^{*})p(\gamma^{*}\mid\bm{G},a,b)}{p(Y\mid X,U,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2})p(\beta,\widetilde{\beta}\mid\gamma,\sigma^{2},h_{\beta})p(\delta,\widetilde{\delta}\mid\gamma)p(\gamma\mid\bm{G},a,b)}$
@@
    $\sigma^{2}\mid y,\bm{X},U,\beta,\widetilde{\beta},\bm{\Omega}\sim IG[a_{\sigma}+\frac{n}{2},b_{\sigma}+\frac{\sum_{i}(y_{i}-x_{i}(\beta+(I-\bm{\Omega}\text{diag}(s))\widetilde{\beta})-U_{i}\widetilde{\beta})^{2}}{2}]$
@@
    $u_{i}\mid y,\bm{X},U,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2}\sim N(\mu_{i},{\bm{\Sigma}}_{U})$
    ${\bm{\Sigma}}_{U}=\left({\bm{A}}+\frac{1}{\sigma^{2}}\widetilde{\beta}\,\widetilde{\beta}^{T}\right)^{-1}$
    $\mu_{i}={\bm{\Sigma}}_{U}\left(\frac{1}{\sigma^{2}}\left(y_{i}\widetilde{\beta}-\widetilde{\beta}\,\beta^{T}x_{i}-\widetilde{\beta}\widetilde{\beta}^{T}\Gamma_{i}\right)\right)$
    $\Gamma=X(I-{\bm{\Omega}}\text{diag}(s)),$